OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). My answer is y = 3x − 14 Am I right? :)

OpenStudy (zzr0ck3r):

3x-y = -3 implies y = 3x+3 so we have slope = 3 plug in your points to check -2 = 3(4) - 14 -2 = -2 all good.

OpenStudy (anonymous):

So its right? :)

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

Ok thank you so much! :)

OpenStudy (zzr0ck3r):

np